# Submitted to Second AEQUITAS Workshop on Fairness and Bias in AI, Co-located with ECAI 2024

### v1

@@ 3.4 Plotting
Correlation- Gender 0.99 $\pm$ .01 0.88 $\pm$ .01 0.86 $\pm$ .01 0.79 $\pm$ .01 0.71 $\pm$ .01 n/a n/a n/a
Remover Race 0.71 $\pm$ .01 0.45 $\pm$ .01 0.86 $\pm$ .01 0.80 $\pm$ .01 0.71 $\pm$ .01 n/a n/a n/a
TabFairGAN Gender 0.69 $\pm$ .01 0.68 $\pm$ .01 0.84$\pm$ .01 0.76 $\pm$ .01 0.65 $\pm$ .01 0.91 $\pm$ .01 0.50 $\pm$ .01 0.537 $\pm$ .01
@@
Threshold Gender 0.93 $\pm$ .01 0.29 $\pm$ .01 0.86 $\pm$ .01 0.75 $\pm$ .01 0.65 $\pm$ .01 n/a n/a n/a
Optimizer Race 0.87 $\pm$ .01 0.40 $\pm$ .01 0.87 $\pm$ .01 0.80 $\pm$ .01 0.71 $\pm$ .01 n/a n/a n/a
@@
Table 4: Evaluation on the Compass dataset using different models presented at the FairX. Bold indicates best result, and all the metrics score are higher as better. Synthetic Data Evaluation is only applicable to the Fair Generative Models (i.e. TabFairGAN and Decaf).
@@
FairDisco Gender 0.99 $\pm$ .01 0.99 $\pm$ .01 0.612 $\pm$ .01 0.61 $\pm$ .01 0.58 $\pm$ .01 n/a n/a n/a
Race 0.99 $\pm$ .01 0.99 $\pm$ .01 0.59 $\pm$ .01 0.59 $\pm$ .01 0.56 $\pm$ .01 n/a n/a n/a

### v2

# Accepted at AEQUITAS 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2023, Santiago de Compostela, Spain

@@ 3.4 Plotting
Correlation- Gender 0.32 $\pm$ .01 0.23 $\pm$ .01 0.86 $\pm$ .01 0.79 $\pm$ .01 0.71 $\pm$ .01 n/a n/a n/a
Remover Race 0.29 $\pm$ .01 0.20 $\pm$ .01 0.86 $\pm$ .01 0.80 $\pm$ .01 0.71 $\pm$ .01 n/a n/a n/a
TabFairGAN Gender 0.69 $\pm$ .01 0.60 $\pm$ .01 0.84$\pm$ .01 0.76 $\pm$ .01 0.65 $\pm$ .01 0.91 $\pm$ .01 0.50 $\pm$ .01 0.537 $\pm$ .01
@@
Threshold Gender 0.95 $\pm$ .01 0.35 $\pm$ .01 0.86 $\pm$ .01 0.66 $\pm$ .01 0.65 $\pm$ .01 n/a n/a n/a
Optimizer Race 0.69 $\pm$ .01 0.25 $\pm$ .01 0.87 $\pm$ .01 0.66 $\pm$ .01 0.71 $\pm$ .01 n/a n/a n/a
@@
Table 4: Evaluation on the Compas dataset using different models presented at the FairX. Bold indicates best result, and all the metrics score are higher as better. Synthetic Data Evaluation is only applicable to the Fair Generative Models (i.e. TabFairGAN and Decaf).
@@
FairDisco Gender 0.97 $\pm$ .01 0.92 $\pm$ .01 0.55 $\pm$ .01 0.54 $\pm$ .01 0.43 $\pm$ .01 n/a n/a n/a
Race 0.87 $\pm$ .01 0.76 $\pm$ .01 0.53 $\pm$ .01 0.53 $\pm$ .01 0.44 $\pm$ .01 n/a n/a n/a
@@ 6 Conclusion and future work
Acknowledgements.
The work was partially funded by the Knut and Alice Wallenberg Foundation, and the TAILOR Network of Excellence for trustworthy AI (EC Grant Agreement 952215). Portions of this work were carried out using the AIOps/Stellar facilities funded by the Excellence Center at Linköping–Lund in Information Technology (ELLIIT).